OpenStudy (anonymous):

Solve the following system of equations. 4x-2y-z=-5 x-3y+2z=3 3x+y-2z=-5

hero (hero):

Learn cramer's rule and you'll have no problem with solving systems such as these

OpenStudy (anonymous):

You can use Cramer's rule, or generate an augmented coefficient matrix and either perform row reduction to get the answer (or the inverse), or use the long, drawn-out analytic way (with the matrix of cofactors and transposes) to get an inverse matrix. Once you have an inverse, you can just multiply it by the right-hand vector, but using Cramer's rule is much simpler.

OpenStudy (anonymous):

What is cramer's rule ?

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Cramer%27s_rule It's important to understand how it works, and once you know, you'll never forget it (it's pretty intuitive). It provides very fast results.

OpenStudy (anonymous):

Go down to "Explicit formulas for small systems".

hero (hero):

For three variable systems, I wouldn't suggest anything else.

OpenStudy (anonymous):

This one's pretty easy though. Wouldn't take more than 3 or 4 minutes to bring it down with row reduction.

hero (hero):

Right, I was assuming the person didn't know how to solve using matrices

OpenStudy (anonymous):

Both are techniques you'd learn in a precalculus or introductory linear algebra course, but you're right - Cramer's rule is pretty much always much simpler and quicker. Less prone to silly arithmetical mistakes, too.

hero (hero):

I could have learned cramers rule when I took algebra for the first time. Not that difficult. Don't know why they avoided teaching it to me when I was a freshie in hs

OpenStudy (anonymous):

There's a lot of baggage associated with linear systems that would logically precede and follow from learning Cramer's rule... it's not difficult for a freshman in high school to learn, but it's just not incredibly relevant (you'd have to invoke matrices, determinants, etc).

hero (hero):

At the very least learning what a 2 x 2 matrix and its determinant would have been a very innocently harmless introduction to what was to come in the future. Better to go into something with one eye rather than with blind eyes.

hero (hero):

0.) Given: 4x-2y-z=-5 x-3y+2z=3 3x+y-2z=-5 1.) Add the 2nd and 3rd Equations together to eliminate z: 4x - 2y = -2 2.) Multiply the first equation by 2 8x - 4y -2z = -10 3.) Add the above result to the second equation to get: 9x - 7y = -7 4.) Use the equations from steps 1.) and 3.) to solve the following system: 4x - 2y = -2 9x - 7y = -7

hero (hero):

I got you started.

OpenStudy (anonymous):

I still don't know . Im just gunna quit and move on .

hero (hero):

I'm sorry to hear that you're going to quit.

hero (hero):

You don't know how to solve system of two equations?

OpenStudy (anonymous):

is it..... (0, 1, 3) ???

hero (hero):

Yes. I'm not going to even ask how you got that

OpenStudy (anonymous):

Hahahahahahaha , okkaaay . :p Thank you though !